OpenStudy (anonymous):

how do i find the equation of the line that intersects the graph of the parabola y=x"2 at x=-2 and x=1?

OpenStudy (amistre64):

by finding the points with the values (-2,(-2)^2) and (1,1^2) and pushing the points thru a line equation setup when you find the slope between them

OpenStudy (anonymous):

can you explain that a little simpler?

OpenStudy (amistre64):

not really

OpenStudy (amistre64):

if you were given 2 points in general, how would you find the equation of the line between them?

OpenStudy (anonymous):

point slopr form

OpenStudy (amistre64):

that would be fine, do you know how to find a slope between 2 points?

OpenStudy (anonymous):

slope=y2-y1/x2-x1

OpenStudy (amistre64):

youre doing great so your real question is what are the points that it is asking for you to use

OpenStudy (anonymous):

well i guess y1 is 4 and x is -2 and then 1 amd 1 ?

OpenStudy (amistre64):

we know that a point is defined as: (x,y) and it is given the y=x^2 which allows us to write the point as: (x,x^2) point1 is when x=-2 point2 is when x =1

OpenStudy (amistre64):

yep, you got them

OpenStudy (amistre64):

(-2,4) and (1,1) are the points they are wanting you to use

OpenStudy (anonymous):

then i use y-y1=m(x-x1) ?

OpenStudy (amistre64):

yep, and the (1,1) looks like an easy point to use in that; what is the slope?

OpenStudy (anonymous):

-1 ?

OpenStudy (amistre64):

correct, m=-1 sooo y-1 = -(x-1) and they might want it in a y= format y = -x+2 seems about right

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

youre welcome :)